OpenStudy (anonymous):

A publisher incurs a fixed expense of $2000 towards composing and other activities in printing a book. A variable cost of $2.25 per book is incurred towards paper, binding charges etc. What is the average cost per book if 40000 copies are printed?

OpenStudy (anonymous):

40000*2.25=90000

ganeshie8 (ganeshie8):

are you saying 90,000 is the cost of one book ? :o

OpenStudy (anonymous):

no .

ganeshie8 (ganeshie8):

good :) keep going..

OpenStudy (anonymous):

40000/2000

ganeshie8 (ganeshie8):

hold up

ganeshie8 (ganeshie8):

total cost for printing 40,000 copies = fixed cost + variable cost

ganeshie8 (ganeshie8):

fixed cost = 2000 variable cost = 2.25*40,000 = 90,000

ganeshie8 (ganeshie8):

total cost = ?

OpenStudy (anonymous):

92000

OpenStudy (anonymous):

so average cost is 920000/40000

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

2.3

OpenStudy (anonymous):

2.30

ganeshie8 (ganeshie8):

looks good ^^